OpenStudy (anonymous):

If the sin of angle x is , and the triangle was dilated to be two times as big as the original, what would be the value of the sin of x for the dilated triangle? Hint: Use the slash symbol ( / ) to represent the fraction bar and enter the fraction with no spaces.

OpenStudy (anonymous):

@jdoe0001 can you help me?

OpenStudy (jdoe0001):

If the sin of angle x is , and the tr ^ ?

OpenStudy (anonymous):

wait what?

OpenStudy (jdoe0001):

well, there seems to be a blank but anyway if the triangle gets dilated, EXPANDED, by twice as much that means the sides get expanded by twice as much too, so 2 * per side so \( sin(\theta)=\cfrac{opposite}{hypotenuse}\quad \begin{array}{llll} 2\cdot opposite\\ 2\cdot hypotenuse \end{array}\implies \bf sin(\theta)=\cfrac{2\cdot opposite}{2\cdot hypotenuse}\)

OpenStudy (anonymous):

so 2/2? or 1

OpenStudy (jdoe0001):

well, depends on what the original value was... you seem to have a blank there

OpenStudy (anonymous):

4/5

OpenStudy (jdoe0001):

If the sin of angle x is , and the tr... ^ blank

OpenStudy (anonymous):

I'm sorry i forgot to add it but its 4/5

OpenStudy (jdoe0001):

\(\bf sin(\theta)=\cfrac{4}{5}\quad \begin{array}{llll} 2\cdot 4\\ 2\cdot 5 \end{array}\implies sin(\theta)=\cfrac{2\cdot 4}{2\cdot 5}\)

OpenStudy (anonymous):

8/10?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (jdoe0001):

ok